Find the class boundaries, midpoints, and size for each class.

(a)  32 – 38

(b)  86 – 104

(c)  895 – 905

(d)  12.3 – 13.5

(e)  3.18 – 4.96 



1
Expert's answer
2022-03-24T19:31:09-0400

a. boundaries:32,38

midpoints: (38+32)/2=35

size: 38-32=6

b.boundaries: 86,104

midpoints: (104+86)/2=95

size: 104-86=18

c. boundaries: 895,905

midpoints: (905+895)/2=900

size: 905-895=10

d. boundaries: 13.5, 12.3

midpoints: (13.5+12.3)/2=12.9

size: 13.5-12.3=1.2

e. boundaries: 3.18, 4.96

midpoints: (3.18+4.96)/2=4.07

size: 4.96-3.18=1,78